OpenStudy (anonymous):

x-coordinate: -41+34-(-18) y-coordinate: -25+56+(-29)

OpenStudy (anonymous):

this is a point (11,2) on graph

OpenStudy (anonymous):

Look : We have : -41+34+18=x -25+56-29=y Now answer again ;)

OpenStudy (anonymous):

x= 52-41 y= 56- 54 Right ?;)

OpenStudy (anonymous):

Do we have to divites

OpenStudy (anonymous):

no. (11,2)